OpenStudy (anonymous):

5(h+2)=-2h(3-5h) solve for specified variable

OpenStudy (ibbutibbu.):

The first step would be to distribute. Do you know how to do that @Felicia123 ?

OpenStudy (anonymous):

yea is that you do the things in ( )

OpenStudy (ibbutibbu.):

No, you multiply everythoing in the parenthesis with whatever is RIGHT outside the parenthesise

OpenStudy (anonymous):

so it will be 5x2= 10 -2x3=-6 -2x-5=10

OpenStudy (ibbutibbu.):

@Felicia123 do you mind me leaving for about 15 minutes? I can help you right after that? is that fine with you?

OpenStudy (anonymous):

yes that is fine

OpenStudy (ibbutibbu.):

Sorry for the trouble, but thank you for letting me go! Will be back in 15!!

OpenStudy (anonymous):

ok lol bye